X has a binomial distribution with number of trials 10 and success probability 0.65. What is the expected value of X?

Group of answer choices

A- 65.00

B- 6.50

C- 0.35

D- 7.00
Answer #1

Solution

Given that ,

n = 10

p = 0.65

Mean = = n * p

= 10 * 0.65 = 6.50

Mean = = 6.50

option B is correct